OpenStudy (anonymous):

What is the solution for x in the proportion 14 over 18 equals x over 63.

OpenStudy (anonymous):

\[\frac{14}{18}=\frac{x}{63}\] cross multiply and get: \[18x=63*14\] \[18x=882\] \[x=\frac{882}{18}\] x=49